OpenStudy (anonymous):

What is the simplified form of the quantity y-squared plus 7y plus 12 over the quantity y-squared minus 3y minus 18 ? Answers: y minus 4 over y minus 6 y minus 4 over y plus 6 y plus 4 over y minus 6 y plus 4 over y plus 6

jimthompson5910 (jim_thompson5910):

\(\Large y^2 + 7y + 12\) factors to what?

OpenStudy (anonymous):

(Y+3)(Y+4)?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

\(\Large y^2 - 3y - 18\) factors to what?

OpenStudy (anonymous):

(y-6)(y+3) not sure about that one

jimthompson5910 (jim_thompson5910):

it's correct

OpenStudy (anonymous):

ok, so what do i do next? cancel out (y+3)?

jimthompson5910 (jim_thompson5910):

Why did we do all that? We factor everything, then we cancel the common (y+3) terms \[\Large \frac{y^2+7y+12}{y^2-3y-18}\] \[\Large \frac{(y+3)(y+4)}{y^2-3y-18}\] \[\Large \frac{(y+3)(y+4)}{(y-6)(y+3)}\] \[\Large \frac{\color{red}{(y+3)}(y+4)}{(y-6)\color{red}{(y+3)}}\] \[\Large \frac{\color{red}{\cancel{(y+3)}}(y+4)}{(y-6)\color{red}{\cancel{(y+3)}}}\] \[\Large \frac{y+4}{y-6}\] ------------------------------------------------------- So \(\Large \frac{y^2+7y+12}{y^2-3y-18}\) simplifies to \(\Large \frac{y+4}{y-6}\)

jimthompson5910 (jim_thompson5910):

yes that's correct

OpenStudy (anonymous):

thank you so much!!!!

jimthompson5910 (jim_thompson5910):

you're welcome